OpenStudy (anonymous):

write an equation in point-slope form for the line through the given point with the given slope.(-3, -5); m = -2/5

OpenStudy (mrnood):

you know the 'point slope' form is y=mx+b and you are told m so the equation is y= -2/5 x +b but you also know that when x= -3 then y = -5 so -5 = -2/5 (-3) + b so oyu can work out b and there is your answer

OpenStudy (er.mohd.amir):

point slope form y-y1=m(x-x1) where m is slope and (x1,y1) is point.

OpenStudy (mathstudent55):

A line with slope m, that passes through and point \((x_1, y_1)\) has the point-slope equation: \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

Thanks! Fastest responses I've ever received lol XD

OpenStudy (mathstudent55):

In your case, you are given \(slope = m = - \dfrac{2}{5} \), and point \((-3, -5)\). Just substitute these values into the equation above. Keep in mind that you need to subtract the x-coordinate and the y-coordinate of the point. Both your x- and y-coordinates are negative, so be careful with the signs.

OpenStudy (mathstudent55):

@MrNood You wrote "you know the 'point slope' form is y=mx+b" ... That is actually the slope-intercept form, not the point-slope form.

OpenStudy (mrnood):

soz - yes . I'll read the question properly next time! :-)

OpenStudy (anonymous):

So, the answer would be y - 5 = -2/5 (x - 3)?

OpenStudy (mathstudent55):

No. That is not the answer. You did plug int he slope correctly, but above I mentioned that the coordinates of the point are both negative numbers and you need to be careful because you are subtracting two negative numbers.

OpenStudy (mathstudent55):

Always start with the equation or formula. Write it down. It will help you with the problem, but it will also help you memorize the formula or equation.

OpenStudy (mathstudent55):

We start with the point-slope equation: \(y-y_1=m(x-x_1) \) Now we replace \(x_1\) with -3, \(y_1\) with -5, and m with \(- \dfrac{2}{5} \) \(y-(-5)=-\dfrac{2}{5}(x-(-3)) \) Notice that all I did was the substitutions for x1, y1 and m. Now simplify the subtractions of negative numbers. Remember that subtracting a negative number is adding. \(y +5=-\dfrac{2}{5}(x+3) \)

OpenStudy (anonymous):

OOoohhhhh. Okie. I see it now...>XD